• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name K. McCormick

    Log entry time 00:14:34 on December18,2003

    Entry number 114557

    keyword=Swing Shift Summary 12/17/03

    ~6:00 pm: Sweep of Hall Commences
    ~7:00 pm: Hall to Beam Permit
    7:30-8:30: Tuning to Hall
    9:00: Ion Chamber Tests Commence
    10:00: I ask for 5 uA to be sent to the Hall, to commence beam

    List of things to do:

    With 5 uA CW, energy and position locks on (FFB not working)
    1. Check BCM Calibration
    Asking for 3 uA CW from the machine, I see .02 uA on the
    Hall A BCMs. The voltage readings from the cavities
    are negligable (.1 mV). However, MCC sees a good spot
    on the dump, so while Doug calls Arun, I'm going to do
    some harp scans. Arun calls, says cavities were retuned
    during down. Says we should call John Musson. I paged
    him. We have a current readback from the MCC screen, so
    we'll use that for the time being.

    2. Harp Scan
    Things aren't working well. It's difficult to get a good
    harp scan, and when we try to analyze the ones we did take,
    the tkl script dies during the analysis. The screen-grabs
    of the error messages are below.

    Doug came in about this time, and went over to the MCC to
    work with them on the spot size.

    FIGURE 1

    FIGURE 2